OpenStudy (anonymous):

come up with vector u and v, with u parallel to y and v perpendicular to y and x= u+v. Y=(cos(pi/6), sin(pi/6)), x= 3(cos(pi/3),sin(pi/3))

OpenStudy (turingtest):

"u parallel to y" means that \[\vec u\cdot\vec y=\|\vec u\|\|\vec y\|\]so take the dot product on the left and multiply the magnitudes right, will get you one equation "v perpendicular to y and x" means that\[\vec u\cdot\vec y=\vec v\cdot\vec x=0\](the dot product of two perpendicular vectors is zero) so do the dot product; that will give you another equation use the system of two equations and solve for the components of \(\vec u\) and \(\vec v\)

OpenStudy (anonymous):

i got u=(\[(-9\sqrt{3}+18)(\sqrt{3}/2,1/2)\] V=(\[16-9\sqrt{3},27-16\sqrt{3}\])

OpenStudy (anonymous):

is that correct?